Company Alpha common stock sells for $39.50 a share and pays an annual dividend that increases by 4 percent annually.

The market rate of return on this stock is 8.45 percent. What is the amount of the last dividend paid?
Answer #1

market rate of return=(D1/Current price)+Growth rate

0.0845=(D1/39.50)+0.04

D1=(0.0845-0.04)*39.50

=$1.75775

Hence last dividend paid=D1*Present value of discounting factor(rate%,time period)

=1.75775/1.04

=$1.69(Approx).
Answer #2

ANSWER :


Given :

r = 8.45 % = 0.0845

P0 = 39.50 ($)

g = 4% = 0.04


Now,

r = D1 / P 0 + g

=> 0.0845 = D1 / 39.50 + 0.04

=> D1 = (0.0845 - 0.04)*39.50 = 1.76 ($) approx.

=> D0 = D1/1.04 = 1.69 ($) approx. 


So,


$1.69   is the last dividend paid  (ANSWER)


answered by: Tulsiram Garg